# Human-AI Collaboration with Misaligned Preferences

### v1

@@ 1 Introduction
In this paper, we focus on a specific instance of human-algorithm collaboration: where the algorithm acts as a curator, and the humans are noisy and potentially misaligned with each other and with the algorithm. Informally, an algorithm acts as a curator when its role is to narrow down a larger set of items to a smaller set, among which the human picks her favorite. This type of role is one of the most common in human-algorithm systems: e.g., in content recommendation, search, and some types of categorical prediction [^1] For example, when a user requests directions, Google maps returns a small set of routes, and the user typically picks her final route from those routes (Figure 1). . A human is noisy if she has inherent randomness in her ability to recognize her true preferences over items. “To err is human”, and it has been widely recognized that humans are often imperfect at picking the “correct” item (e.g., see (Agranov and Ortoleva, 2017; Plackett, 1975; Luce et al., 1959; Hey and Orme, 1994)). A human is misaligned with another agent (human or algorithm) if they have different ground-truth preferences over items, which is distinct from their potentially noisy realizations of those preferences. For example, if Alice prefers horror movies and Bob prefers comedies, then their ground truth preferences over a set of movies would be misaligned with each other. Additionally, given a single algorithmic recommender, either Alice or Bob (or both) would be misaligned with the algorithm. Of course, in the limiting case of perfect personalization, both could have separate algorithms that are perfectly aligned with their preferences, but in realistic settings where one algorithm must serve a diverse range of humans, misalignment is a reality of life. Note that our framework has strong connections to other bodies of work, such as pluralistic alignment and conformal prediction: we discuss such connections in Section 3.
@@
  • An algorithm that maximizes utilitarian welfare is always noiseless. While it is in general computationally hard to find this algorithm, we can design a mixed integer program (MIP) which is similar to the MIP used for assortment optimization under the Mallows model (Désir et al., 2016).
@@ 2.1 Human-algorithm Collboration Model
In a human-algorithm collaboration model, there are two actors: an algorithm ($A$) and a human ($H$). There is a set of items $M=\{x_{1},\ldots,x_{m}\}$ representing different outcomes (e.g., labels in prediction tasks, generative model output). The setting we study is that of where the algorithm acts as a curator, picking a subset of size $k$ from a larger set of items of size $m$. The specific mode of interaction we assume is where the algorithm picks their top $k$ from some noisy ranking over items, and the human picks their favorite among that set, according to their own noisy ranking over items [^2] Note that this is the same modeling assumption as (Donahue et al., 2024) in the unanchored case, but our results significantly generalize over their results by allowing agents to be misaligned. This setting may occur in settings such as content recommendation (e.g., Netflix or Spotify) or algorithmic assistants in predictive tasks.
@@ 2.2 Noisy Ranking Distribution
We assume that each agent can only access a noisy permutation over items: this may be because the agent only has access to imperfect or noisy data about each alternative, for example, or due to features such as temperature in generative AI tools (Ackley et al., 1985). Denote the sampled rankings of the algorithm and human as $\operatorname{\pi}\sim\mathcal{D}_{a}$ and $\operatorname{\rho}_{i}\sim\mathcal{D}_{h}^{i}$ respectively, where $\mathcal{D}_{a}$ concentrates at ground-truth ranking $\operatorname{\pi}^{*}$ and $\mathcal{D}_{h}^{i}$ concentrates at $\operatorname{\rho}_{i}^{*}$. Denote by $x_{i}\succ_{\pi}x_{j}$ that $x_{i}$ precedes $x_{j}$ in $\pi$. Denote by $\pi(i)$ and $\pi[:k]$ the $i$-th item and the first $k$ items (viewed as an unordered set) in a ranking $\operatorname{\pi}$. Denote by $\pi\circ(x_{i}x_{j})$ the ranking after swapping the locations of $x_{i}$ and $x_{j}$ in ranking $\pi$. All distributions in the paper are inversion-monotonic and label-invariant. Informally, adding inversion to a ranking only decreases its probability, and the distribution is not tied to the labels of the items. Both Mallows model (Mallows, 1957) and Plackett-Luce model (Luce et al., 1959; Plackett, 1975) satisfy the two properties.
@@
The Mallows model (Mallows, 1957) generates a distribution over permutations based on their number of inversions relative to a central ranking (also known as the Kendall-Tau distance). A Mallows model $\mathcal{D}(\pi^{*},\phi)$ consists of a central ranking $\pi^{*}$ and an accuracy parameter $\phi\geq 0$. The probability of a permutation $\pi$ occurring is $\operatorname*{\mathbb{P}}[\pi]=\frac{1}{Z}\cdot\exp(-\phi\cdot d(\pi^{*},\pi))$, where $Z$ is the normalization constant and $d(\pi^{*},\pi)$ is the Kendall-Tau distance between $\pi$ and $\pi^{*}$. We assume all accuracy parameters used in this paper are positive.
@@
The random utility model (RUM) is commonly used as a model of permutations (Thurstone, 1927), where a permutation is generated as follows: i.i.d. noise is added to the true value of each item to produce $\hat{v_{i}}=v_{i}+\epsilon$, for $\epsilon\sim\mathcal{D}$: then, the items are sorted in order of the noised values $\{\hat{v_{i}}\}$. Note that in a RUM, the probability of sampling a particular ranking depends not only on the number of inversions but also on the specific valuations of the items. Although there is no definition of “central ranking” in RUM, we refer to the ranking consistent with the item values as the central ranking. When $\epsilon$ follows i.i.d. Gumbel noise $G(\mu,\beta)$, following the analysis in (Luce et al., 1959; Plackett, 1975), the probability of a permutation $\sigma$ is given by
@@ 2.3 Welfare objectives
A human-algorithm joint system achieves uplift if the expected utility of every type of human is improved: $\operatorname*{\mathbb{E}}[u_{i}(x_{C}^{i})]>\operatorname*{\mathbb{E}}[u_{i}(x_{H}^{i})]$ for any $i\in[n]$ [^3] We refer the reader to the related work section for discussion about the relation between uplift and complementarity (Bansal et al., 2021b).
@@ 3 Related Work
Our work relates to the general area of human-algorithm collaboration. In particular, there is a rich history of applied and empirical work in human-algorithm interaction: we refer interested readers to see (Preece et al., 1994; Kim, 2015; MacKenzie, 2024; Lazar et al., 2017) for textbook treatments. Specifically, our work relates to a growing literature using theoretical models to analyze humans interacting with algorithms. Some works study how to design algorithms to optimally assist humans (Bansal et al., 2021a; Chan et al., 2019; Donahue et al., 2022; Madras et al., 2018; Brown and Agarwal, 2022), including work incorporating models of human cognitive biases, such as (Li et al., 2024; Ibrahim et al., 2025; Chen et al., 2025). Other work decides when human-algorithm teams perform well (Greenwood et al., 2024; Peng et al., 2024; Steyvers et al., 2022; Brown and Agarwal, 2024; Cowgill and Stevenson, 2020; Green and Chen, 2019; Alur et al., 2023, 2024; Guo et al., 2025), often relating to benchmarks such as complementarity (strict improvement over the human or algorithm alone defined in (Bansal et al., 2021b)). Some relevant literature reviews, taxonomies, and systematic studies in this space include (Gomez et al., 2025; Vaccaro et al., 2024; Rastogi et al., 2022).
@@
Within human-algorithm collaboration, our work is most closely related to that of Donahue et al. (2024), which studies a similar setting where an algorithmic tool presents a top $k$ subset to a noisy human. A key difference in our work is that we allow humans to be misaligned with the algorithm and study when misalignment is helpful. In fact, many of our contributions strictly generalize theirs, such as generalizing the utility function beyond top item recovery. Other related areas include conformal prediction, which studies how to optimize a subset of items (e.g., ensuring that the best item is presented with high probability): see (Angelopoulos et al., 2023; Fontana et al., 2023) for a summary of work in this area. Within this space, some works focus on optimizing the set of items that are presented (Straitouri et al., 2022; Wang et al., 2022; Straitouri and Rodriguez, 2023; De Toni et al., 2024; Babbar et al., 2022; De Toni et al., 2024; Hullman et al., 2025; Zhang et al., 2024), while others include more empirical analyses of specific settings (Angelopoulos et al., 2020; Arnaiz-Rodriguez et al., 2025). In general, these works do not consider settings with multiple humans who may be misaligned with each other: one exception is (Corvelo Benz and Gomez Rodriguez, 2025), which studies an empirical setting on when algorithmic alignment is a helpful property for tools assisting human decision-makers. There is also a line of work studying human-AI collaboration through information elicitation (Steyvers et al., 2022; Corvelo Benz and Rodriguez, 2023; Corvelo Benz and Gomez Rodriguez, 2025; Collina et al., 2024, 2025a). Steyvers et al. (2022) developed a Bayesian framework to combine the predictions of humans and AI algorithms with confidence scores. Corvelo Benz and Rodriguez (2023); Corvelo Benz and Gomez Rodriguez (2025) then showed that if the confidence value aligns with the human’s confidence, the collaboration will benefit the human’s decision-making. Collina et al. (2024, 2025a) studied the setting where each party holds different feature information and transmits their numerical prediction to each other at each round. Collina et al. (2024) generalizes Aumann’s Agreement Theorem by relaxing the rationality assumptions and introducing calibration-based conditions on each party that ensure efficient convergence of the conversation. Building on this, Collina et al. (2025a) then proposes more communication-efficient protocols and removes the assumption of a common and correct prior. Finally, Collina et al. (2025b) models a Bayesian communication setting where a human is relying on information from multiple AI tools that are misaligned with each other (and with her) to make a discrete decision. Compared to their work, our work does not model the interaction of the human and the algorithm as a multi-round process. The algorithm acts more like a curator, instead of a rational agent.
@@
Another research area that our paper relates to is pluralistic alignment. In this setting, the goal is generally to align an algorithm tool with users who have heterogenous preferences (e.g. (Sorensen et al., 2024)), especially the work connected with voting and social choice literature on aggregating diverse user preferences (e.g. (Conitzer et al., 2024; Dai and Fleisig, 2024; Shirali et al., 2025; A Alamdari et al., 2024; Ge et al., 2024; Pardeshi et al., 2024; Chen et al., 2024; Siththaranjan et al., 2023; Gölz et al., 2025)). One key difference between much of this work and our own is the meaning of noise. Often, work in this space (and voting in general) assumes that users know their own preferences and can represent them faithfully, which in the case of political opinions is often a fairly reasonable assumption. In our setting, we are assuming that humans may be able to only noisily access their true values over items and thus can benefit from the assistance of another agent (e.g., an algorithm). This is a more natural assumption in lower-stakes settings such as content recommendation or many types of content creation, such as writing code or daily emails. Separately, some papers study the performance of voting rules under noise (e.g., sporting competition), though they tend to be less focused on societal welfare objectives (Boehmer et al., 2022; Caragiannis et al., 2016).
@@
Note that the term of uplift is related to that of complementarity as defined in (Bansal et al., 2021b): “a human-AI system achieves complementary performance if it outperforms both the AI and the human acting alone”. As compared to complementarity, uplift is weaker because it only requires that the human-algorithm system outperform the human alone, rather than the human and algorithm jointly. However, it also differs (and is weakly stronger) in that it is a population-level property: an algorithm must benefit all humans using it in order to achieve uplift. In addition, the application scenarios of complementarity and uplift also differ. The classical setting of complementarity usually assumes that both the human and the algorithm aim to predict an objective event. In contrast, our model does not assume the existence of an objective ranking of the items, and the goal of human-algorithm collaboration is to better align humans’ preferences. Therefore, the performance of the algorithm alone is not well-defined in our setting and depends on how well the algorithm meets the human’s preferences.
@@
The basic ranking model considered in our paper is defined over the number of inversions. A wide range of alternative permutation models has been studied in the literature. For example, the Bradley-Terry model (Bradley and Terry, 1952) and the Plackett-Luce model (Plackett, 1975; Luce et al., 1959) view a permutation as the consequence of a sequence of choices; the weighted Mallows model (Raman and Joachims, 2014) assigns weights to the inversions and defines the probability of a permutation by weighted Kendall-Tau distance. Moreover, Awasthi et al. (2014); Liu and Moitra (2018) studied learning a mixture of the Mallows model, which assumes the sampled permutation comes from a heterogeneous population.
@@ 4.2 Extensions to Specific Distributions
Sufficient condition for when collaboration with a misaligned algorithm is helpful for the human under the Plackett-Luce model is when some $i^{\prime}\in[i-1]$ satisfies
@@ 5.2 Uplift
Note that Lemma˜6 strictly generalizes results in (Donahue et al., 2024), which provided these results only for when exactly $2$ items are presented and the human is in the top item recovery case (only has value for her top item). Moreover, Lemma˜6 does not put any constraint on the zero-valued items, and the algorithm and human can be misaligned on these items. The proof can be found at Section˜C.3.
@@ 5.2.2 Small Noise Facilitates Uplift
We remark that Lemma˜7’s message is in line with other works on fairness (e.g., (Jain et al., 2024; Singh and Joachims, 2018)) that deem randomization to be helpful for fairness. One difference in our setting is that we expect that in many cases there will be relatively large benefits in terms of uplift from relatively small levels of noise (and thus, relatively small impact on the utilitarian welfare objective). This is because small amounts of noise that result in occasionally including a new item has asymmetric costs and benefits: it has large benefits for users who stronger prefer that item to those that are already included because they will be able to select it with relatively high probability. However, it has small negative impacts on users who have low value for that item, since on the occasions when it is included, they will have a high probability of correctly ignoring that item as irrelevant.

### v2

@@ 1 Introduction
In this paper, we focus on a specific instance of human-algorithm collaboration: where the algorithm acts as a curator, and the humans are noisy and potentially misaligned with each other and with the algorithm. Informally, an algorithm acts as a curator when its role is to narrow down a larger set of items to a smaller set, among which the human picks her favorite. This type of role is one of the most common in human-algorithm systems: e.g., in content recommendation, search, and some types of categorical prediction [^1] For example, when a user requests directions, Google maps returns a small set of routes, and the user typically picks her final route from those routes (Figure 1). . A human is noisy if she has inherent randomness in her ability to recognize her true preferences over items. “To err is human”, and it has been widely recognized that humans are often imperfect at picking the “correct” item (e.g., see (agranov2017stochastic; plackett1975analysis; luce1959individual; hey1994investigating)). A human is misaligned with another agent (human or algorithm) if they have different ground-truth preferences over items, which is distinct from their potentially noisy realizations of those preferences. For example, if Alice prefers horror movies and Bob prefers comedies, then their ground truth preferences over a set of movies would be misaligned with each other. Additionally, given a single algorithmic recommender, either Alice or Bob (or both) would be misaligned with the algorithm. Of course, in the limiting case of perfect personalization, both could have separate algorithms that are perfectly aligned with their preferences, but in realistic settings where one algorithm must serve a diverse range of humans, misalignment is a reality of life. Note that our framework has strong connections to other bodies of work, such as pluralistic alignment and conformal prediction: we discuss such connections in Section 3.
@@
  • An algorithm that maximizes utilitarian welfare is always noiseless. While it is in general computationally hard to find this algorithm, we can design a mixed integer program (MIP) which is similar to the MIP used for assortment optimization under the Mallows model (desir2016assortment).
@@ 2.1 Human-algorithm Collboration Model
In a human-algorithm collaboration model, there are two actors: an algorithm ($A$) and a human ($H$). There is a set of items $M=\{x_{1},\ldots,x_{m}\}$ representing different outcomes (e.g., labels in prediction tasks, generative model output). The setting we study is that of where the algorithm acts as a curator, picking a subset of size $k$ from a larger set of items of size $m$. The specific mode of interaction we assume is where the algorithm picks their top $k$ from some noisy ranking over items, and the human picks their favorite among that set, according to their own noisy ranking over items [^2] Note that this is the same modeling assumption as (donahue2024listsbetteronebenefits) in the unanchored case, but our results significantly generalize over their results by allowing agents to be misaligned. This setting may occur in settings such as content recommendation (e.g., Netflix or Spotify) or algorithmic assistants in predictive tasks.
@@ 2.2 Noisy Ranking Distribution
We assume that each agent can only access a noisy permutation over items: this may be because the agent only has access to imperfect or noisy data about each alternative, for example, or due to features such as temperature in generative AI tools (ackley1985learning). Denote the sampled rankings of the algorithm and human as $\operatorname{\pi}\sim\mathcal{D}_{a}$ and $\operatorname{\rho}_{i}\sim\mathcal{D}_{h}^{i}$ respectively, where $\mathcal{D}_{a}$ concentrates at ground-truth ranking $\operatorname{\pi}^{*}$ and $\mathcal{D}_{h}^{i}$ concentrates at $\operatorname{\rho}_{i}^{*}$. Denote by $x_{i}\succ_{\pi}x_{j}$ that $x_{i}$ precedes $x_{j}$ in $\pi$. Denote by $\pi(i)$ and $\pi[:k]$ the $i$-th item and the first $k$ items (viewed as an unordered set) in a ranking $\operatorname{\pi}$. Denote by $\pi\circ(x_{i}x_{j})$ the ranking after swapping the locations of $x_{i}$ and $x_{j}$ in ranking $\pi$. All distributions in the paper are inversion-monotonic and label-invariant. Informally, adding inversion to a ranking only decreases its probability, and the distribution is not tied to the labels of the items. Both Mallows model (mallows1957non) and Plackett-Luce model (luce1959individual; plackett1975analysis) satisfy the two properties.
@@
The Mallows model (mallows1957non) generates a distribution over permutations based on their number of inversions relative to a central ranking (also known as the Kendall-Tau distance). A Mallows model $\mathcal{D}(\pi^{*},\phi)$ consists of a central ranking $\pi^{*}$ and an accuracy parameter $\phi\geq 0$. The probability of a permutation $\pi$ occurring is $\operatorname*{\mathbb{P}}[\pi]=\frac{1}{Z}\cdot\exp(-\phi\cdot d(\pi^{*},\pi))$, where $Z$ is the normalization constant and $d(\pi^{*},\pi)$ is the Kendall-Tau distance between $\pi$ and $\pi^{*}$. We assume all accuracy parameters used in this paper are positive.
@@
The random utility model (RUM) is commonly used as a model of permutations (thurstone1927law), where a permutation is generated as follows: i.i.d. noise is added to the true value of each item to produce $\hat{v_{i}}=v_{i}+\epsilon$, for $\epsilon\sim\mathcal{D}$: then, the items are sorted in order of the noised values $\{\hat{v_{i}}\}$. Note that in a RUM, the probability of sampling a particular ranking depends not only on the number of inversions but also on the specific valuations of the items. Although there is no definition of “central ranking” in RUM, we refer to the ranking consistent with the item values as the central ranking. When $\epsilon$ follows i.i.d. Gumbel noise $G(\mu,\beta)$, following the analysis in (luce1959individual; plackett1975analysis), the probability of a permutation $\sigma$ is given by
@@ 2.3 Welfare objectives
A human-algorithm joint system achieves uplift if the expected utility of every type of human is improved: $\operatorname*{\mathbb{E}}[u_{i}(x_{C}^{i})]>\operatorname*{\mathbb{E}}[u_{i}(x_{H}^{i})]$ for any $i\in[n]$ [^3] We refer the reader to the related work section for discussion about the relation between uplift and complementarity (bansal2021does).
@@ 3 Related Work
Our work relates to the general area of human-algorithm collaboration. In particular, there is a rich history of applied and empirical work in human-algorithm interaction: we refer interested readers to see (preece1994human; kim2015human; mackenzie2024human; lazar2017research) for textbook treatments. Specifically, our work relates to a growing literature using theoretical models to analyze humans interacting with algorithms. Some works study how to design algorithms to optimally assist humans (bansal2021most; chan2019assistive; donahue2022human; madras2017predict; Agarwal2022DiversifiedRF), including work incorporating models of human cognitive biases, such as (li2024decoding; ibrahim2025measuring; chen2025missing). Other work decides when human-algorithm teams perform well (greenwood2024designing; peng2024no; steyvers2022bayesian; agarwal2023online; cowgill2020algorithmic; benCSCW; alur2023auditing; alur2024human; guo2025value), often relating to benchmarks such as complementarity (strict improvement over the human or algorithm alone defined in (bansal2021does)). Some relevant literature reviews, taxonomies, and systematic studies in this space include (gomez2025human; vaccaro2024combinations; rastogi2022unifying).
@@
Within human-algorithm collaboration, our work is most closely related to that of donahue2024listsbetteronebenefits, which studies a similar setting where an algorithmic tool presents a top $k$ subset to a noisy human. A key difference in our work is that we allow humans to be misaligned with the algorithm and study when misalignment is helpful. In fact, many of our contributions strictly generalize theirs, such as generalizing the utility function beyond top item recovery. Other related areas include conformal prediction, which studies how to optimize a subset of items (e.g., ensuring that the best item is presented with high probability): see (angelopoulos2023conformal; fontana2023conformal) for a summary of work in this area. Within this space, some works focus on optimizing the set of items that are presented (straitouri2022provably; wang2022improving; straitouri2023designing; de2024towards; babbar2022utility; de2024towards; hullman2025conformal; zhang2024evaluating), while others include more empirical analyses of specific settings (angelopoulos2020uncertainty; arnaiz2025towards). In general, these works do not consider settings with multiple humans who may be misaligned with each other: one exception is (corvelo2025human), which studies an empirical setting on when algorithmic alignment is a helpful property for tools assisting human decision-makers. There is also a line of work studying human-AI collaboration through information elicitation (steyvers2022bayesian; corvelo2023human; corvelo2025human; collina2024tractable; collina2025collaborative). steyvers2022bayesian developed a Bayesian framework to combine the predictions of humans and AI algorithms with confidence scores. corvelo2023human; corvelo2025human then showed that if the confidence value aligns with the human’s confidence, the collaboration will benefit the human’s decision-making. collina2024tractable; collina2025collaborative studied the setting where each party holds different feature information and transmits their numerical prediction to each other at each round. collina2024tractable generalizes Aumann’s Agreement Theorem by relaxing the rationality assumptions and introducing calibration-based conditions on each party that ensure efficient convergence of the conversation. Building on this, collina2025collaborative then proposes more communication-efficient protocols and removes the assumption of a common and correct prior. Finally, collina2025emergent models a Bayesian communication setting where a human is relying on information from multiple AI tools that are misaligned with each other (and with her) to make a discrete decision. Compared to their work, our work does not model the interaction of the human and the algorithm as a multi-round process. The algorithm acts more like a curator, instead of a rational agent.
@@
Another research area that our paper relates to is pluralistic alignment. In this setting, the goal is generally to align an algorithm tool with users who have heterogenous preferences (e.g. (sorensen2024roadmap)), especially the work connected with voting and social choice literature on aggregating diverse user preferences (e.g. (conitzer2024social; dai2024mapping; shirali2025direct; a2024policy; ge2024axioms; pardeshi2024learning; chen2024pal; siththaranjan2023distributional; golz2025distortion)). One key difference between much of this work and our own is the meaning of noise. Often, work in this space (and voting in general) assumes that users know their own preferences and can represent them faithfully, which in the case of political opinions is often a fairly reasonable assumption. In our setting, we are assuming that humans may be able to only noisily access their true values over items and thus can benefit from the assistance of another agent (e.g., an algorithm). This is a more natural assumption in lower-stakes settings such as content recommendation or many types of content creation, such as writing code or daily emails. Separately, some papers study the performance of voting rules under noise (e.g., sporting competition), though they tend to be less focused on societal welfare objectives (boehmer2022quantitative; 10.1145/2892565).
@@
Note that the term of uplift is related to that of complementarity as defined in (bansal2021does): “a human-AI system achieves complementary performance if it outperforms both the AI and the human acting alone”. As compared to complementarity, uplift is weaker because it only requires that the human-algorithm system outperform the human alone, rather than the human and algorithm jointly. However, it also differs (and is weakly stronger) in that it is a population-level property: an algorithm must benefit all humans using it in order to achieve uplift. In addition, the application scenarios of complementarity and uplift also differ. The classical setting of complementarity usually assumes that both the human and the algorithm aim to predict an objective event. In contrast, our model does not assume the existence of an objective ranking of the items, and the goal of human-algorithm collaboration is to better align humans’ preferences. Therefore, the performance of the algorithm alone is not well-defined in our setting and depends on how well the algorithm meets the human’s preferences.
@@
The basic ranking model considered in our paper is defined over the number of inversions. A wide range of alternative permutation models has been studied in the literature. For example, the Bradley-Terry model (bradley1952rank) and the Plackett-Luce model (plackett1975analysis; luce1959individual) view a permutation as the consequence of a sequence of choices; the weighted Mallows model (raman2014methods) assigns weights to the inversions and defines the probability of a permutation by weighted Kendall-Tau distance. Moreover, awasthi2014learning; liu2018efficiently studied learning a mixture of the Mallows model, which assumes the sampled permutation comes from a heterogeneous population.
@@ 4.2 Extensions to Specific Distributions
Sufficient condition for when collaboration with a misaligned algorithm is helpful under the Plackett-Luce model for the human is when some $i^{\prime}\in[i-1]$ satisfies
@@ 5.2 Uplift
Note that Lemma˜6 strictly generalizes results in (donahue2024listsbetteronebenefits), which provided these results only for when exactly $2$ items are presented and the human is in the top item recovery case (only has value for her top item). Moreover, Lemma˜6 does not put any constraint on the zero-valued items, and the algorithm and human can be misaligned on these items. The proof can be found at Section˜C.3.
@@ 5.2.2 Small Noise Facilitates Uplift
We remark that Lemma˜7’s message is in line with other works on fairness (e.g., (jain2024scarce; singh2018fairness)) that deem randomization to be helpful for fairness. One difference in our setting is that we expect that in many cases there will be relatively large benefits in terms of uplift from relatively small levels of noise (and thus, relatively small impact on the utilitarian welfare objective). This is because small amounts of noise that result in occasionally including a new item has asymmetric costs and benefits: it has large benefits for users who stronger prefer that item to those that are already included because they will be able to select it with relatively high probability. However, it has small negative impacts on users who have low value for that item, since on the occasions when it is included, they will have a high probability of correctly ignoring that item as irrelevant.